# Exploiting Agent Symmetries for Performance Analysis of Distributed Optimization Methods

### v1

@@ 2.2.2 A relaxed formulation
where $X_{\mathbin{\!/\mkern-5.0mu/\!}},Y_{\mathbin{\!/\mkern-5.0mu/\!}}\in\mathbb{R}^{nd\times t}$ contains columns of $X$ and $Y$ projected onto the subspace $\mathcal{B}$, and $X_{\perp},Y_{\perp}\in\mathbb{R}^{nd\times t}$ the columns projected on its orthogonal complement $\mathcal{B}^{\perp}$.
Démonstration.
We need to prove that conditions (50), (51) and (52) are necessary and sufficient for the existence of a matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{B}}$ such that $Y=MX$. The proof relies on the use of [1, Theorem 3.3] which proved the necessity and sufficiency of interpolation constraints similar to (51) and (52), for general linear operators, without any preserved subspace. In our case, we know that the matrix $M$ has no impact on the subspace $\mathcal{B}$ and only acts on the part in $\mathcal{B}^{\perp}$. We thus need to use [1, Theorem 3.3] on this part only. Decoupling both parts can be done using an appropriate change of variables:
@@
where $Q=[Q_{\mathbin{\!/\mkern-5.0mu/\!}}\leavevmode\nobreak\ Q_{\perp}]\in\mathbb{R}^{nd\times nd}$ is an orthogonal matrix of change of variables and is split into two sets of columns: $Q_{\mathbin{\!/\mkern-5.0mu/\!}}\in\mathbb{R}^{nd\times|\mathcal{B}|}$ form a basis of the subspace $\mathcal{B}$ and $Q_{\perp}\in\mathbb{R}^{nd\times|\mathcal{B}^{\perp}|}$ a basis of $\mathcal{B}^{\perp}$. The new variable $\tilde{X}$ has thus two sets of components, $\tilde{X}_{\mathbin{\!/\mkern-5.0mu/\!}}\in\mathbb{R}^{|\mathcal{B}|\times t}$ describing the coordinates along the subspace $\mathcal{B}$, and $\tilde{X}_{\perp}\in\mathbb{R}^{|\mathcal{B}^{\perp}|\times t}$ describing the coordinates along the orthogonal complement $\mathcal{B}^{\perp}$. By definition, $X_{\mathbin{\!/\mkern-5.0mu/\!}}$ and $Y_{\mathbin{\!/\mkern-5.0mu/\!}}$ have components only along $\mathcal{B}$ and $X_{\perp}$ and $Y_{\perp}$ only along $\mathcal{B}^{\perp}$:
$X_{\mathbin{\!/\mkern-5.0mu/\!}}$ $=Q\begin{bmatrix}\tilde{X}_{\mathbin{\!/\mkern-5.0mu/\!}}\\ 0\end{bmatrix},\qquad Y_{\mathbin{\!/\mkern-5.0mu/\!}}=Q\begin{bmatrix}\tilde{Y}_{\mathbin{\!/\mkern-5.0mu/\!}}\\ 0\end{bmatrix},$ and $X_{\perp}=Q\begin{bmatrix}0\\ \tilde{X}_{\perp}\end{bmatrix},\qquad Y_{\perp}=Q\begin{bmatrix}0\\ \tilde{Y}_{\perp}\end{bmatrix},$ (62)
By applying this change of variable to condition (50), we obtain $\tilde{X}_{\mathbin{\!/\mkern-5.0mu/\!}}=\tilde{Y}_{\mathbin{\!/\mkern-5.0mu/\!}}$. The two conditions are well equivalent because the change of variable is revertible. By applying this change of variable (62) to conditions (51) and (52), since $Q$ is orthogonal, i.e. $Q^{T}Q=I$, we obtain
$(\tilde{Y}_{\perp}-\lambda^{-}\tilde{X}_{\perp})^{T}(\tilde{Y}_{\perp}-\lambda^{+}\tilde{X}_{\perp})$ $\preceq 0,$ (63)
@@
As the change of variable is revertible, the two sets of constraints are equivalent. By [1, Theorem 3.3], these two constraints are necessary and sufficient for the existence of a symmetric matrix $\tilde{M}_{\perp}\in\mathbb{R}^{|\mathcal{B}^{\perp}|\times|\mathcal{B}^{\perp}|}$ with eigenvalues between $\lambda^{-}$ and $\lambda^{+}$ such that $\tilde{Y}_{\perp}=\tilde{M}_{\perp}\tilde{X}_{\perp}$. Therefore, we have proved that conditions (50), (51) and (52) are equivalent to
$\begin{bmatrix}\tilde{Y}_{\mathbin{\!/\mkern-5.0mu/\!}}\\ \tilde{Y}_{\perp}\end{bmatrix}=\begin{bmatrix}I_{d}&0\\ 0&\tilde{M}_{\perp}\end{bmatrix}\begin{bmatrix}\tilde{X}_{\mathbin{\!/\mkern-5.0mu/\!}}\\ \tilde{X}_{\perp}\end{bmatrix}\qquad\text{ with $\tilde{M}_{\perp}$ symmetric and $\lambda(\tilde{M}_{\perp})\in[\lambda^{-},\lambda^{+}]$}.$ (65)
We now need to come back to the initial variables $X$ and $Y$ and characterize the matrix $M$ in terms of $\tilde{M}_{\perp}$. We can use the (revertible) change of variable (53) to express equation (65) in terms of $Y$ and $X$
$Y=Q\begin{bmatrix}I_{d}&0\\ 0&\tilde{M}_{\perp}\end{bmatrix}Q^{T}X\qquad\text{ with $\lambda(\tilde{M}_{\perp})\in[\lambda^{-},\lambda^{+}]$}.$ (66)
By splitting $Q$ as $Q=[Q_{\mathbin{\!/\mkern-5.0mu/\!}}\leavevmode\nobreak\ Q_{\perp}]$, matrix $M$ can be written as
@@
Since $\tilde{M}_{\perp}$ is symmetric, this expression shows that $M$ is also symmetric. We now show that symmetric matrix $M$ has $|\mathcal{B}|$ eigenvalues equal to $1$, with associated eigenvectors $Q_{\mathbin{\!/\mkern-5.0mu/\!}}$. Let $v_{\mathbin{\!/\mkern-5.0mu/\!}}$ denote any column of $Q_{\mathbin{\!/\mkern-5.0mu/\!}}$, i.e. any basis vector of the subspace $\mathcal{B}$, we have well that
@@
The other eigenvalues of $M$ are equal to the ones of $\tilde{M}_{\perp}$. Indeed, Let $(v_{\perp},\lambda_{\perp})$ be a pair of eigenvector and eigenvalue of matrix $\tilde{M}_{\perp}\in\mathbb{R}^{\mathcal{B}^{\perp}\times\mathcal{B}^{\perp}}$, then $(Q_{\perp}v_{\perp},\lambda_{\perp})$ is a pair of eigenvector and eigenvalue of matrix $M\in\mathbb{R}^{nd\times nd}$. This can be verified easily as
@@
In Theorem 4, we have proved that conditions (50), (51) and (52) are necessary and sufficient for the existence of a matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{B}}$ such that $Y=MX$. In the case of consensus steps (6), we can apply Theorem 4 with $\mathcal{B}$ being the consensus subspace $\mathcal{C}$. Indeed, since $\quantity(\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}\otimes I_{d})\subset\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{C}}$, any set of pairs of points which is $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable is also $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{C}}$-interpolable. The reverse statement does not hold since a matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{C}}$ can apply different weights to each agent component and may also mix different components between them, which is not the case for consensus steps of the form (6). This means that Theorem 4 with $\mathcal{B}=\mathcal{C}$ provides non-tight interpolation constraints for $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, leading to potentially non-tight performance bounds when exploited in PEP for distributed optimization. One can always check the tightness of the PEP solution a posteriori, by checking if the provided solution is $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable. If successful, this check also recovers an instance of the worst matrix $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, see [5] for details.
@@
While the literature for distributed optimization usually assumes consensus steps of the form $\mathbf{y}=(W\otimes I_{d})\mathbf{x}$ with $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, the existing theoretical performance guarantees also appear to use non-tight description of these consensus steps. Indeed, we did not find any theoretical guarantee that exploits the shape $(W\otimes I_{d})$ of the consensus steps. In particular, many results remain valid for generalized steps $\mathbf{y}=M\mathbf{x}$ with $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{C}}$ because many proofs rely on the convergence analysis of the consensus step, which is not impacted by the shape of $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{C}}$, as shown in Proposition 5. This proposition extends classical results of consensus theory to matrices $M\in\mathbb{R}^{nd\times nd}$ that do not have the shape of $(W\otimes I_{d})$.
@@
[…]
@@
Proof of Proposition 5.
Since $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{C}}$, the matrix admits the following spectral decomposition:
$M=V\Sigma V^{T},\qquad\text{with}\qquad V_{1,\dots,d}=(\mathbf{1}_{n}\otimes I_{d})/{\scriptstyle\sqrt{n}},$
@@
$\mathbf{z}^{k+1}=\Sigma\mathbf{z}^{k}$
@@
where vector $\tilde{\mathbf{z}}^{0}$ contains the $d$ first entries of $\mathbf{z}^{0}$ and $(n-1)d$ zeros. Thus, the sequence $\mathbf{x}^{k}$ converges to $V\tilde{\mathbf{z}}^{0}$ which is the combination of the columns $V_{1,\dots,d}$ with coefficients given by the $d$ first entries of $\mathbf{z}^{0}$, i.e.
$\lim_{k\to\infty}\mathbf{x}^{k}=\lim_{k\to\infty}V\mathbf{z}^{k}=V\tilde{\mathbf{z}}^{0}=V_{1,\dots,d}\leavevmode\nobreak\ \mathbf{z}^{0}_{1,\dots,d}.$
@@
$\lim_{k\to\infty}\mathbf{x}^{k}=V_{1,\dots,d}V_{1,\dots,d}^{T}\mathbf{x}^{0}=\frac{1}{n}(\mathbf{1}_{n}\mathbf{1}_{n}^{T}\otimes I_{d})\mathbf{x}^{0}=\mathbf{1}_{n}\otimes\overline{x}^{0}.$
One can show that the interpolation constraints stated in Theorem 4 are LMI Gram-representable when $\mathcal{B}=\mathcal{C}$ and by Proposition 2, they can therefore be used in SDP PEP formulations. The main idea is that the constraints can all be expressed using only the scalar products of the local variables $x_{i}^{k},y_{i}^{k}$. The detailed proof is given in [5, Theorem 1] but we can observe that all projection matrices can be expressed based on the agents variables matrices: the projections $X_{\mathbin{\!/\mkern-5.0mu/\!}},\leavevmode\nobreak\ Y_{\mathbin{\!/\mkern-5.0mu/\!}}\in\mathbb{R}^{nd\times t}$ on the consensus subspace, can be computed with the agent average of $X$ and $Y$: $X_{\mathbin{\!/\mkern-5.0mu/\!}}=\mathbf{1}_{n}\otimes\overline{X}$ and $Y_{\mathbin{\!/\mkern-5.0mu/\!}}=\mathbf{1}_{n}\otimes\overline{Y}$, with $\overline{X}^{k}=\frac{1}{n}\sum_{i=1}^{n}x_{i}^{k}$, and the projections on the orthogonal complement can be computed as $X_{\perp}=X-X_{\mathbin{\!/\mkern-5.0mu/\!}}$ and $Y_{\perp}=X-Y_{\mathbin{\!/\mkern-5.0mu/\!}}$. These interpolation constraints can thus be used to define a convex set of Gram matrices, which contains the non-convex set $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$.

### v2

@@ 2.2.2 A relaxed formulation
where $X_{\mathbin{\!/\mkern-5.0mu/\!}},Y_{\mathbin{\!/\mkern-5.0mu/\!}}\in\mathbb{R}^{nd\times t}$ contains columns of $X$ and $Y$ projected onto the subspace $\mathcal{C}$, and $X_{\perp},Y_{\perp}\in\mathbb{R}^{nd\times t}$ the columns projected on its orthogonal complement $\mathcal{C}^{\perp}$.
Remark.
Theorem 4 is stated for the case where symmetric linear operators from $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ preserve the consensus subspace $\mathcal{C}$ (Definition 6), but it actually holds for any general preserved subspace $\mathcal{B}\subseteq\mathbb{R}^{nd}$. Indeed, the proof below remains valid if we replace the consensus subspace $\mathcal{C}$ by any subspace $\mathcal{B}$ (and adapt the dimensions where needed).
Proof of Theorem 4.
We need to prove that conditions (48), (49) and (50) are necessary and sufficient for the existence of a matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ such that $Y=MX$. The proof relies on the use of [1, Theorem 3.3] which proved the necessity and sufficiency of interpolation constraints similar to (49) and (50), for general linear operators, without any preserved subspace. In our case, we know that the matrix $M$ has no impact on the consensus subspace $\mathcal{C}$ and only acts on the part in $\mathcal{C}^{\perp}$. We thus need to use [1, Theorem 3.3] on this part only. Decoupling both parts can be done using an appropriate change of variables :
@@
where $Q=[Q_{\mathbin{\!/\mkern-5.0mu/\!}}\penalty 10000\ Q_{\perp}]\in\mathbb{R}^{nd\times nd}$ is an orthogonal matrix of change of variables and is split into two sets of columns : $Q_{\mathbin{\!/\mkern-5.0mu/\!}}\in\mathbb{R}^{nd\times|\mathcal{C}|}$ form a basis of the consensus subspace $\mathcal{C}$ and $Q_{\perp}\in\mathbb{R}^{nd\times|\mathcal{C}^{\perp}|}$ a basis of its orthogonal complement $\mathcal{C}^{\perp}$. The new variable $\tilde{X}$ has thus two sets of components, $\tilde{X}_{\mathbin{\!/\mkern-5.0mu/\!}}\in\mathbb{R}^{|\mathcal{C}|\times t}$ describing the coordinates along the subspace $\mathcal{C}$, and $\tilde{X}_{\perp}\in\mathbb{R}^{|\mathcal{C}^{\perp}|\times t}$ describing the coordinates along the orthogonal complement $\mathcal{C}^{\perp}$. By definition, $X_{\mathbin{\!/\mkern-5.0mu/\!}}$ and $Y_{\mathbin{\!/\mkern-5.0mu/\!}}$ have components only along $\mathcal{C}$ and $X_{\perp}$ and $Y_{\perp}$ only along $\mathcal{C}^{\perp}$ :
$X_{\mathbin{\!/\mkern-5.0mu/\!}}$ $=Q\begin{bmatrix}\tilde{X}_{\mathbin{\!/\mkern-5.0mu/\!}}\\ 0\end{bmatrix}$, $Y_{\mathbin{\!/\mkern-5.0mu/\!}}$ $=Q\begin{bmatrix}\tilde{Y}_{\mathbin{\!/\mkern-5.0mu/\!}}\\ 0\end{bmatrix}$, and $X_{\perp}$ $=Q\begin{bmatrix}0\\ \tilde{X}_{\perp}\end{bmatrix}$, $Y_{\perp}$ $=Q\begin{bmatrix}0\\ \tilde{Y}_{\perp}\end{bmatrix}$. (53)
By applying this change of variable to condition (48), we obtain $\tilde{X}_{\mathbin{\!/\mkern-5.0mu/\!}}=\tilde{Y}_{\mathbin{\!/\mkern-5.0mu/\!}}$. The two conditions are well equivalent because the change of variable is reversible. By applying this change of variable (LABEL:eq:change_of_var_split) to conditions (49) and (50), since $Q$ is orthogonal, i.e., $Q^{T}Q=I$, we obtain
[…]
@@
As the change of variable is reversible, the two sets of constraints are equivalent. By [1, Theorem 3.3], these two constraints are necessary and sufficient for the existence of a symmetric matrix $\tilde{M}_{\perp}\in\mathbb{R}^{|\mathcal{C}^{\perp}|\times|\mathcal{C}^{\perp}|}$ with eigenvalues between $\lambda^{-}$ and $\lambda^{+}$ such that $\tilde{Y}_{\perp}=\tilde{M}_{\perp}\tilde{X}_{\perp}$. Therefore, we have proved that conditions (48), (49) and (50) are equivalent to
$\begin{bmatrix}\tilde{Y}_{\mathbin{\!/\mkern-5.0mu/\!}}\\ \tilde{Y}_{\perp}\end{bmatrix}=\begin{bmatrix}I_{d}&0\\ 0&\tilde{M}_{\perp}\end{bmatrix}\begin{bmatrix}\tilde{X}_{\mathbin{\!/\mkern-5.0mu/\!}}\\ \tilde{X}_{\perp}\end{bmatrix}\penalty 10000\ \penalty 10000\ \text{with $\tilde{M}_{\perp}=\tilde{M}_{\perp}^{T}$,\penalty 10000\ \penalty 10000\ $\lambda(\tilde{M}_{\perp})\in[\lambda^{-},\lambda^{+}]$}.$ (58)
We now need to come back to the initial variables $X$ and $Y$ and characterize the matrix $M$ in terms of $\tilde{M}_{\perp}$. We can use the (reversible) change of variable (51) to express equation (58) in terms of $Y$ and $X$
$Y=\underbrace{Q\begin{bmatrix}I_{d}&0\\ 0&\tilde{M}_{\perp}\end{bmatrix}Q^{T}}_{M}X\qquad\text{ with $\lambda(\tilde{M}_{\perp})\in[\lambda^{-},\lambda^{+}]$}.$ (59)
By splitting $Q$ as $Q=[Q_{\mathbin{\!/\mkern-5.0mu/\!}}\penalty 10000\ Q_{\perp}]$, matrix $M$ can be written as
@@
Since $\tilde{M}_{\perp}$ is symmetric, this expression shows that $M$ is also symmetric. We now show that the symmetric matrix $M$ has $|\mathcal{C}|$ eigenvalues equal to $1$, with associated eigenvectors $Q_{\mathbin{\!/\mkern-5.0mu/\!}}$. Let $v_{\mathbin{\!/\mkern-5.0mu/\!}}$ denote any column of $Q_{\mathbin{\!/\mkern-5.0mu/\!}}$, i.e., any basis vector of the consensus subspace $\mathcal{C}$, we have well that
@@
The other eigenvalues of $M$ are equal to the ones of $\tilde{M}_{\perp}$. Indeed, Let $(v_{\perp},\lambda_{\perp})$ be a pair of eigenvector and eigenvalue of matrix $\tilde{M}_{\perp}\in\mathbb{R}^{\mathcal{C}^{\perp}\times\mathcal{C}^{\perp}}$, then $(Q_{\perp}v_{\perp},\lambda_{\perp})$ is a pair of eigenvector and eigenvalue of matrix $M\in\mathbb{R}^{nd\times nd}$. This can be verified easily as
@@
In Theorem 4, we have proved that conditions (48), (49) and (50) are necessary and sufficient for the existence of a matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ such that $Y=MX$. We now give some interpretation of the constraints, explain why they are only necessary for $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolability, and describe the impacts in the PEPs for decentralized optimization.
Interpretation of the constraints
By Definition 6 of the consensus subspace, the orthogonal decomposition terms can be expressed using the agent averages :
$\hskip 56.9055ptX_{\mathbin{\!/\mkern-5.0mu/\!}}$ $=\overline{X}\otimes\mathbf{1}_{n}$ $X_{\perp}$ $=X-X_{\mathbin{\!/\mkern-5.0mu/\!}},$ (63)
$\hskip 56.9055ptY_{\mathbin{\!/\mkern-5.0mu/\!}}$ $=\overline{Y}\otimes\mathbf{1}_{n}$ $Y_{\perp}$ $=X-Y_{\mathbin{\!/\mkern-5.0mu/\!}},$ (64)
where $\overline{X},\penalty 10000\ \overline{Y}\in\mathbb{R}^{d\times t}$ are agent average of $X$ and $Y$, i.e., $\overline{X}_{\cdot,k}=\frac{1}{n}\sum_{i=1}^{n}x_{i}^{k}$. Therefore, constraint (48) is equivalent to $\overline{X}=\overline{Y}$. This constraint is related to the $1$ eigenvalues of the matrix and their corresponding eigenvectors, which form a basis for the consensus space $\mathcal{C}$. The constraint imposes that each vector $\mathbf{x}^{k}$ have the same agents average as $\mathbf{y}^{k}$, for each generalized consensus step (38) ($k=1,\dots,t$).
Interpretations of the linear matrix inequality (LMI) constraints (49) and (50) are detailed in [5]. In summary, (49) requires the disagreement between the agents, measured by $y_{\perp}^{T}y_{\perp}$ for $y$ and $x_{\perp}^{T}x_{\perp}$ for $x$ to be reduced by a factor $\lambda^{2}\in[0,1)$ after each consensus step. Constraints (49) and (50) also allow linking different consensus steps to each other, via the impact of off-diagonal terms which exploit the fact that these steps use the same averaging matrix.
Necessary interpolation conditions for $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$
In Theorem 4, we have proved that conditions (48), (49) and (50) are necessary and sufficient for the existence of a matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ such that $Y=MX$. By definition of $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ and $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, these conditions are necessary for the existence of an averaging matrix $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ such that $Y=(W\otimes I_{d})X$, as shown in the following corollary.
Corollary 4.1.
Let $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ be a set of $t$ pairs of points and $\lambda^{-},\lambda^{+}\in(-1,1)$, $\lambda^{-}\leq\lambda^{+}$. If $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ is $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable (Definition 4), then we have
$\overline{X}$ $=\overline{Y},$ (65)
$X_{\perp}^{T}Y_{\perp}$ $=Y_{\perp}^{T}X_{\perp},$ (67)
with $\overline{X},\overline{Y}\in\mathbb{R}^{d\times t}$ and $X_{\perp},Y_{\perp}\in\mathbb{R}^{nd\times t}$ defined in (63)-(64).
Démonstration.
From a matrix $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, we can always build a matrix $M=(W\otimes I_{d})\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, meaning that any set of pairs of points that is $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable is also $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable. Therefore, the points $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ are $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable and satisfy constraints (65), (66) and (67) by Theorem 4. ∎
The reverse statement does not hold since constraints (65), (66) and (67) guarantee the interpolation by a matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, that may not have the Kronecker structure required to extract an averaging matrix $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$. The full mask of $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ allows different weights to be applied to each component of the agent and can also mix different components together, which is not the case for consensus steps (21) that impose a sparse mask of the form $(W\otimes I_{d})$.
Relaxed Performance Estimations Problems
One can show that the interpolation constraints stated in Corollary 4.1 are LMI Gram-representable. The detailed proof is given in [5, Theorem 1] but the main idea is to express the constraints with the scalar products of the local variables $x_{i}^{k},y_{i}^{k}$. Therefore, by Proposition 2, these constraints can be used in SDP PEP formulations. As explained in [5], using necessary constraints (65), (66) and (67) to represent consensus steps of the form of (21) in PEP allows obtaining worst-case performance guarantees for a decentralized optimization algorithm $\mathcal{A}\in\mathcal{A}_{D}$, that are valid for any averaging matrix $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$. This formulation is therefore called the spectral agent-dependent PEP formulation and its results the spectral worst-case. This is an agent-dependent PEP formulation because the PEP depends on each agent individually.
When different averaging matrices are used for different sets of consensus steps in the algorithm, these constraints from Corollary 4.1 can be applied independently to each set of consensus steps.
The bounds obtained with this formulation have a priori no tightness guarantees because it considers larger sets of possible consensus iterates $\{\mathbf{x}^{k},\mathbf{y}^{k}\}$, i.e., those that are $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable. However, this relaxation of the Kronecker structure in consensus steps does not appear to impact the worst-case results, as already observed numerically in [5]. Moreover, one can always check the tightness of the PEP solution a posteriori, by checking if the provided solution is $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable. If successful, this check also recovers an instance of the worst matrix $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, see [5] for details.
Impact of the relaxation
@@
While the literature for distributed optimization usually assumes consensus steps of the form $\mathbf{y}=(W\otimes I_{d})\mathbf{x}$ with $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, the existing theoretical performance guarantees also appear to use non-tight description of these consensus steps. Indeed, we did not find any theoretical guarantee that exploits the Kronecker structure $(W\otimes I_{d})$ of the consensus steps. In particular, many results in decentralized optimization remain valid for generalized consensus steps $\mathbf{y}=M\mathbf{x}$ with $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ because the proofs generally rely on the convergence analysis of the consensus step, which is not impacted by the structure of $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, as shown in Proposition 5. This proposition extends classical results of consensus theory, to the use of matrices $M\in\mathbb{R}^{nd\times nd}$ that do not have the Kronecker structure of $(W\otimes I_{d})$. When $d=1$, we have $M=W\in\mathbb{R}^{n\times n}$ and no extension is needed.
@@
Let $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}\subset\mathbb{R}^{nd\times nd}$ with $\lambda^{-},\lambda^{+}\in(-1,1)$ and $\mathbf{x}^{0}\in\mathbb{R}^{nd}$ a vector stacking the variables $x_{i}^{0}\in\mathbb{R}^{d}$ of each agent. If the vector $\mathbf{x}^{0}$ is updated with a series of general matrix steps $\mathbf{x}^{k+1}=M\mathbf{x}^{k}$, then the sequence of vectors $\mathbf{x}^{k}$ converges to the agent average consensus vector, i.e., a vector stacking $n$ times the agent average $\overline{x}^{0}=\frac{1}{n}\sum_{i=1}^{n}x_{i}^{0}$ :
@@
Démonstration.
Since $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, the matrix admits the following spectral decomposition :
$M=V\Sigma V^{T},\qquad\text{with}\qquad V_{1,\dots,d}=(\mathbf{1}_{n}\otimes I_{d})/{\scriptstyle\sqrt{n}},$ (68)
@@
$\mathbf{z}^{k+1}=\Sigma\mathbf{z}^{k}.$
@@
where vector $\tilde{\mathbf{z}}^{0}$ contains the $d$ first entries of $\mathbf{z}^{0}$ and (n1)d(n-1)d zeros. Thus, the sequence $\mathbf{x}^{k}$ converges to $V\tilde{\mathbf{z}}^{0}$ which is the combination of the columns $V_{1,\dots,d}$ with coefficients given by the $d$ first entries of $\mathbf{z}^{0}$, i.e.,
$\lim_{k\to\infty}\mathbf{x}^{k}=\lim_{k\to\infty}V\mathbf{z}^{k}=V\tilde{\mathbf{z}}^{0}=V_{1,\dots,d}\penalty 10000\ \mathbf{z}^{0}_{1,\dots,d}.$
@@
$\lim_{k\to\infty}\mathbf{x}^{k}=V_{1,\dots,d}V_{1,\dots,d}^{T}\mathbf{x}^{0}=\frac{1}{n}(\mathbf{1}_{n}\mathbf{1}_{n}^{T}\otimes I_{d})\mathbf{x}^{0}=\mathbf{1}_{n}\otimes\overline{x}^{0},$
where the second equality holds by definition of $V_{1,\dots,d}$ (68). ∎
2.3 Section summary : agent-dependent PEP